OpenStudy (anonymous):

Let a = b a^2 = ab a^2-b^2 = ab - b^2 (a+b)(a-b) = b(a-b) a+b = b a+a = a 2a = a 2 = 1 Explain.

OpenStudy (anonymous):

You're dividing by zero on step 3... \[(a+b)(a-b)=b(a-b)\\a+b=b\] The above algebraic transformation is only well-defined for where...\[a-b\neq0\]